Car of mass 200kg travels around a fat circular racetrack of radius 85m. The car starts at resy and its speed increases at constant rate, tangential acceleration:0.6ms^-2. What is the speed of the car when its centripetal and tangential acceleration are equal?
1
Expert's answer
2012-08-14T10:02:24-0400
at = 0.6 m/s^2. Since at = ar,
ar = V^2 / r
V = sqrt(ar * r) = sqrt(at *
r)
V = sqrt(0.6 * 85) = 7.14 m/s